Find the convergence of the following series,

  1. βˆ‘βˆž 𝑛=1(βˆšπ‘›+1βˆ’βˆšπ‘›)/ 𝑛 Β 
  2. βˆ‘βˆž 𝑛=1 cos(1/n)
  3. βˆ‘βˆž 𝑛=1 (sin n)/n
  4. βˆ‘βˆž 𝑛=1 ((𝑛!) 2)/((2𝑛)!)
  5. βˆ‘βˆž 𝑛=1 (n/n+1)n2
1
Expert's answer
2021-11-08T02:49:49-0500

1.

"\displaystyle\sum_{n=1}^{\infin}\dfrac{\sqrt{n+1}-\sqrt{n}}{n}"

"=\displaystyle\sum_{n=1}^{\infin}\dfrac{\sqrt{n+1}-\sqrt{n}}{n}(\dfrac{\sqrt{n+1}+\sqrt{n}}{\sqrt{n+1}+\sqrt{n}})"

"=\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n(\sqrt{n+1}+\sqrt{n})}"

"=\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n^{3/2}(\sqrt{1+1/n}+1)}"

The p-series "\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n^{3/2}}" is convergent (p=3/2).


"\dfrac{1}{n^{3/2}(\sqrt{1+1/n}+1)}<\dfrac{1}{2n^{3/2}}, n\geq1"

Then the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n^{3/2}(\sqrt{1+1/n}+1)}" converges by the Comparison Test.


Therefore the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{\sqrt{n+1}-\sqrt{n}}{n}" converges by the Comparison Test.

2.

"\displaystyle\sum_{n=1}^{\infin}\cos(\dfrac{1}{n})"

"\lim\limits_{n\to\infin}\cos(\dfrac{1}{n})=\cos(0)=1\not=0"

Therefore the series "\displaystyle\sum_{n=1}^{\infin}\cos(\dfrac{1}{n})" diverges by the Test for Divergence.


3.

"\displaystyle\sum_{n=1}^{\infin}\dfrac{\sin(n)}{n}"

Use the Dirichlet's Test


"2\sin(1)\displaystyle\sum_{n=1}^{m}\sin(n)=\displaystyle\sum_{n=1}^{m}(\cos(n-1)-\cos(n+1))"

"=(\cos(0)-\cos(2))+(\cos(1)-\cos(3))"

"+(\cos(2)-\cos(4))+...+(\cos(m-2)-\cos(m))"

"+(\cos(m-1)-\cos(m+1))"

"=1+\cos(1)-\cos(m)-\cos(m+1)"

Then


"\displaystyle\sum_{n=1}^{m}\sin(n)=\dfrac{1+\cos(1)-\cos(m)-\cos(m+1)}{2\sin(1)}"

"\bigg|\displaystyle\sum_{n=1}^{m}\sin(n)\bigg|=\dfrac{|1+\cos(1)-\cos(m)-\cos(m+1)|}{2\sin(1)}"

"\leq \dfrac{1+1+1+1}{2\sin(1)}=\dfrac{2}{\sin(1)}=M"

Let "a_n=\dfrac{1}{n}." Then

"a_{n+1}=\dfrac{1}{n+1}<\dfrac{1}{n}=a_n, n\geq1""\lim\limits_{n\to\infin}a_n=\lim\limits_{n\to\infin}\dfrac{1}{n}=0"

Therefore the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{\sin(n)}{n}" converges by the Dirichlet's Test.


4.

"\displaystyle\sum_{n=1}^{\infin}\dfrac{(n!)^2}{(2n)!}"

Apply the Ratio Test.


"\lim\limits_{n\to\infin}\dfrac{\dfrac{((n+1)!)^2}{(2(n+1))!}}{\dfrac{(n!)^2}{(2n)!}}=\lim\limits_{n\to\infin}\dfrac{(n+1)^2}{(2n+1)(2n+2)}=\dfrac{1}{4}<1"

Therefore the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{(n!)^2}{(2n)!}" converges by the RatioTest.


5.

"\displaystyle\sum_{n=1}^{\infin}(\dfrac{n}{n+1})^{n^2}"

Use the Root Test.


"\\lim\\limits_{n\\to\\infin}\\sqrt[n]{(\\dfrac{n}{n+1})^{n^2}}=\\lim\\limits_{n\\to\\infin}(1-\\dfrac{1}{n}\t\n)^n=e^{-1}=\\dfrac{1}{e}<1"

Therefore the series "\displaystyle\sum_{n=1}^{\infin}(\dfrac{n}{n+1})^{n^2}" converges by the Root Test.